OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -36 and 2304, respectively

OpenStudy (binary3i):

if you divide 5th term with 2ed you'll get cube of the common ratio. once you get the common ratio then square it and divide 2ed term which is -36 you'll get the scale factor.

OpenStudy (anonymous):

Could you walk me through it? This makes no sense to me

OpenStudy (binary3i):

ok. see this is GP \[a,\ ar,\ ar^2,\ ar^3,\ ar^4,\ \ldots\] second term is \[ar\]=-36 fifth term is \[a*r^4=\] 2304 fifth term divided with second term ar^4/ar =r^3={2304/(-36)}=-64 r=-4 now divide second term which is ar=-36 with r ar/r = a =-36/(-4) = 9 a=9 so nth term is \[a*r^{n-1}\] where a is 9 and r is -4

OpenStudy (anonymous):

an = so 9 x 4^(n-1)

OpenStudy (binary3i):

its -4 not 4

OpenStudy (anonymous):

Than you so much!!